OpenStudy (henrietepurina):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)=(x+a)/b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (henrietepurina):

do you think you can help?

OpenStudy (henrietepurina):

@hero ? Please :)

hero (hero):

\(f(x) = \dfrac{x + a}{b}\) \(g(x) = cx - d\)

OpenStudy (henrietepurina):

yup thats it :)

hero (hero):

And in order for \(f(x)\) and \(g(x)\) to be inverses: \(f(g(x)) = x\) and \(g(f(x)) = x\)

OpenStudy (henrietepurina):

ok....

hero (hero):

What if a = d b = c

OpenStudy (henrietepurina):

that cant be possible because a can only equal a for example 1 cant be 2, 1 is 1

hero (hero):

Basically, from the work I just did, the only way for f(x) and g(x) to be inverses, that is the condition that must be met a = d b = c

hero (hero):

So for example, if \(f(x) = \dfrac{x + 3}{4}\) then \(g(x) = 4x - 3\)

OpenStudy (henrietepurina):

aaaaahhh.... now i get it

hero (hero):

Because \(f(g(x)) = \dfrac{(4x - 3) + 3}{4} = \dfrac{4x}{4} =x \)

OpenStudy (henrietepurina):

so is that the answer for part 1?

OpenStudy (henrietepurina):

oh, ok so what about part 2?

hero (hero):

\(g(f(x)) = 4\left(\dfrac{x + 3}{4}\right) - 3 = x\)

OpenStudy (henrietepurina):

is that also the work?

hero (hero):

Actually, I didn't notice all the different parts that you have.

OpenStudy (henrietepurina):

ohhh, sorry :(

hero (hero):

Okay, so to find the inverse of f(x) for example: \(f(x) = \dfrac{x + 3}{4}\)

hero (hero):

Let y = f(x): \(y = \dfrac{x + 3}{4}\)

hero (hero):

Then swap x and y: \(x = \dfrac{y + 3}{4}\)

hero (hero):

And isolate y again: \(4x = y + 3\) \(4x - 3 = y\) Then replace y with \(g(x)\) \(4x - 3 = g(x)\)

hero (hero):

For part 4, you can use desmos.com for plotting graphs.

OpenStudy (henrietepurina):

yeah , i use it all the time :)

OpenStudy (henrietepurina):

Could you rewrite part 3?

hero (hero):

\(f(g(x)) = \dfrac{(4x - 3) + 3}{4} = \dfrac{4x}{4} =x\) \(g(f(x)) = 4\left(\dfrac{x + 3}{4}\right) - 3 = x\)

OpenStudy (henrietepurina):

And then 2?

hero (hero):

I just went over part 2 above. It starts with: "to find the inverse of f(x)"

OpenStudy (henrietepurina):

so wait is part 1 the same as 3?

hero (hero):

Part 1 is to find values for a, b, c, and d.

OpenStudy (henrietepurina):

AHH... OK!

OpenStudy (henrietepurina):

what is a, b, c, and d?

hero (hero):

This is the point where you should be able to think for yourself and go over this information on your own to figure out the values of a, b, c, d. Start from the very beginning and read over this discussion and try to figure out the values on your own.

OpenStudy (henrietepurina):

is a and d 3 and b and d 4?

hero (hero):

where's c?

OpenStudy (henrietepurina):

the d in the first one was supposed to be a c

hero (hero):

Re-type the whole thing you are proposing here.

OpenStudy (henrietepurina):

is a and c 4? and is b and d 3?

OpenStudy (henrietepurina):

ok i got it: a - 3 b - 4 c - 4 d - 3

OpenStudy (henrietepurina):

@Hero